OpenStudy (anonymous):

find the soloution of the system: y= x^2 - 2x - 2 y= 4x + 5

Parth (parthkohli):

2x = x^2 - y - 2.... SO 4x = 2x^2 - 2y - 4 y = (2x^2 - 2y - 4) + 5 y = 2x^2 - 2y - 4 + 5 3y = 2x^2 + 1 CAN YOU SOLVE NOW SUBSTITUTING x?

OpenStudy (anonymous):

i still need help, i dont understand.

OpenStudy (anonymous):

X=7 Y=33 X=-1 Y=1

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

x=7 y=33 x=-1 y=1